OpenStudy (anonymous):

How do I graph y = x(x - 4)

mathslover (mathslover):

first find x intercept then y intercept

OpenStudy (anonymous):

x = 4, y = -4

mathslover (mathslover):

put y = 0 and then solve for x what do you get @deehthebee ?

OpenStudy (campbell_st):

well the parabola is concave down and has x intercepts at 0 and 4 the line of symmetry is x = 2 and the minimum value occurs when x = 2... substitute to find that value. find the y- intercept by letting x = 0...

mathslover (mathslover):

y=x(x-4) put y = 0 0 = x(x-4) x = 0 or x = 4 put x = 0 now y= 0

OpenStudy (anonymous):

What do you mean?

OpenStudy (anonymous):

@mathslover

mathslover (mathslover):

what I meant is what campbell explained

OpenStudy (anonymous):

I don't really get it. Can you explain it again please?

mathslover (mathslover):

oh k no problem i will do

mathslover (mathslover):

y = x(x-4) put y = 0 first.. solve for x what do you get?

OpenStudy (anonymous):

y = 0, x = 4

mathslover (mathslover):

right can x also be 0 ?

OpenStudy (anonymous):

I don't think so?

mathslover (mathslover):

why 0 = x(x-4) divide both sides by x-4 0/(x-4) = x(x-4)/(x-4) 0 = x

OpenStudy (anonymous):

Oh I see. So how would I draw this then?

mathslover (mathslover):

first of all the line of symetry x = 2 u can continue now

OpenStudy (anonymous):

I've never heard of that. What do you mean?

OpenStudy (campbell_st):

every parabola is symmetrical...

OpenStudy (anonymous):

But are you saying that it has to go through x = 2?

OpenStudy (anonymous):

@mathslover

OpenStudy (anonymous):

Hello?

OpenStudy (lgbasallote):

@mathslover i think someone's calling you for the last quarter of an hour

OpenStudy (anonymous):

Yes I have lol

OpenStudy (lgbasallote):

maybe @mathslover didnt know the next step so he quit? :p haha lol

OpenStudy (anonymous):

Strange because mathslover has viewed the thread a few times but never responded

OpenStudy (anonymous):

Yay I am getting a response!

OpenStudy (lgbasallote):

seems my reverse psych worked

OpenStudy (anonymous):

Well, I was getting a response...

mathslover (mathslover):

I am typing wait

OpenStudy (anonymous):

OK :)

mathslover (mathslover):

sorry for giving late response but wait for my response

mathslover (mathslover):

mathslover is not typing a reply I am typing it in notepad will copy paste it .. please don't think -ve

OpenStudy (lgbasallote):

my work here is done

mathslover (mathslover):

here i go